SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (April 18, 2006)

Attendees: Rainer Bartoldus, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su

Operations

  • PEP is probably going to have a two-day downtime later next week to replace a Q5 chamber due to vacuum problems. This would be a good opportunity to do some Trigger work if any.
  • Chris O'Grady posted a HN which discussed the deadtime problem.

DCT/DCZ

  • We continue to have problem configure DCZ due to one of the TSF boards not detected at configureation when user reset was done. This is associated with the firmware issue. Here is a log showing how frequent this has happened over the last two weeks or so. If there is no progress in solving the DCM locking problem, we will roll back to the previous version of the TSF firmware during the next downtime.
  • The DCC board at the teststand is one that just got repaired from France, and the flaky one will be shipped to France for repair.
  • Fang is looking at the headers of some out-of-order events that were flagged as DAQ damaged by L3.

GLT

  • Su Dong is working on the trigger configuratons for the B factory operation report, and will provide options for what we can do in terms of trigger rate.
  • Rainer made a trigger configuration named "phyiscs_tight" which tightens up some of the trigger lines. This will be tested at IR2 soon.
  • There were some glitches that caused GLT damages seen in run 63834. See HN posted.

L3

  • Steffen Luitz and Rainer set up a farm node to test L3 which shows roughly a factor of two for the new box (on one half of its two dual-core CPUs). We're planning to buy ~50 of these boxes by this summer for L3 use.

Data Quality and Monitoring

  • Rahmat showed us some strip charts for the ratio of lumi's. Gaussian errors were supposed to be used in these plots, but it was suspected there are still problems in getting the correct error bars for these plots.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.